OpenStudy (anonymous):

what is log x-log 2=1

Vocaloid (vocaloid):

basic log rules log(A)-log(B) = log(A/B) log(x)-log(2) = log(x/2) log(x/2) = 1

Vocaloid (vocaloid):



Vocaloid (vocaloid):



OpenStudy (anonymous):

thank you once again